A company is planning to spend $16,115 for a share repurchase program. The company's current EPS is $1.33 per share. The current share price is $53.70 per share, and there are 2,684 common shares currently outstanding. Ignoring taxes, what is the company's P/E ratio immediately after the share repurchase has been completed?
Answer #1

Net income = Number of shares * EPS

Net income = 2,684 * 1.33 = 3,569.72

Number of shares repurchased = 16,115 / 53.7 = 300

Outstanding shares after repurchasing = 2,684 - 300

Outstanding shares = 2,384

EPS = Net income / outstanding shares

EPS = 3,569.72 / 2,384

EPS = 1.497366

PE = Price / EPS

PE = 53.7 / 1.497366

PE = 35.86